OpenStudy (anonymous):

Math Check help please! I need to turn one equation into y=mx+b (slope intercept form) and I wanted someone to check me. The equation is: 9x=8y+5 I got: y= -9/8y + -5/8 Also my school website does not except mixed numbers, so all fractions are improper fractions. Any help is appreciated.

OpenStudy (anonymous):

9x=8y+5 9x-5=8y (9/8)x-5/8=y

OpenStudy (anonymous):

so its y=9/8x - 5/8?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you very much. I was darn close which is comforting.

OpenStudy (anonymous):

If you need it in improper fractrion form

OpenStudy (anonymous):

9/8=\[1\frac{1}{8}\] You can get that by dividing 9 by 8

OpenStudy (anonymous):

yea my school website refuses to take mixed numbers so no 1 1/8 :(